OpenStudy (anonymous):

consider the equation y = √x Find dx/dt, given x = 64 and dy/dt = 7.

OpenStudy (tkhunny):

We have y = y(t). Then y' = dy/dt We also have x = x(t). What is \(\dfrac{d}{dt}\sqrt{x}\)? You'll need a good Chain Rule.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

A solution using Mathematica is attached.